### v1

# NoMIRACL: Knowing When You Don’t Know for Robust Multilingual Retrieval-Augmented Generation

Abstract
Retrieval-augmented generation (RAG) grounds large language model (LLM) output by leveraging external knowledge sources to reduce factual hallucinations. However, prior works lack a comprehensive evaluation of different language families, making it challenging to evaluate LLM robustness against errors in external retrieved knowledge. To overcome this, we establish NoMIRACL, a human-annotated dataset for evaluating LLM robustness in RAG across 18 typologically diverse languages. NoMIRACL includes both a non-relevant and a relevant subset. Queries in the non-relevant subset contain passages manually judged as non-relevant or noisy, whereas queries in the relevant subset include at least a single judged relevant passage. We measure LLM robustness using two metrics: (i) hallucination rate, measuring model tendency to hallucinate an answer, when the answer is not present in passages in the non-relevant subset, and (ii) error rate, measuring model inaccuracy to recognize relevant passages in the relevant subset. We build a GPT-4 baseline which achieves a 33.2% hallucination rate on the non-relevant and a 14.9% error rate on the relevant subset on average. Our evaluation reveals that GPT-4 hallucinates frequently in high-resource languages, such as French or English. This work highlights an important avenue for future research to improve LLM robustness to learn how to better reject non-relevant information in RAG. [^1] NoMIRACL dataset and the GPT-4 baseline are available at: https://github.com/project-miracl/nomiracl
1 Introduction
Retrieval Augmented Generation (RAG) (Guu et al., 2020; Lewis et al., 2020; Izacard and Grave, 2021; Borgeaud et al., 2022) is a promising way to incorporate external knowledge via a first-stage retrieval system. RAG instills information from reliable knowledge corpora (provided as external passages) to generate accurate and faithful responses Shuster et al. (2021); Gao et al. (2023b); Liu et al. (2023). Ever since the introduction of large language models (LLMs), such as GPT-3 Brown et al. (2020) or LLAMA-2 Touvron et al. (2023), they are the de-facto choice for generation in RAG, due to their impressive performance across multiple reasoning and natural language inference tasks Chang et al. (2023); Guo et al. (2023). RAG grounds the LLM-generated output, thereby avoiding previously observed factual hallucination Maynez et al. (2020); Raunak et al. (2021) and outdated knowledge Cao et al. (2021); He et al. (2023) in LLMs.
Figure 1: LLM robustness evaluation as a binary tree in NoMIRACL. When dealing with queries in the non-relevant subset ($\mathrm{F}$), the LLM is expected to disregard all noisy passages and refrain from answering ($\mathrm{N}$). Conversely, for queries in the relevant subset ($\mathrm{T}$), the LLM should recognize the relevant passage and provide a valid answer ($\mathrm{P}$).
A challenging issue in RAG is to provide robust and reliable LLM-generated output. The generation stage is dependent on the first-stage information retrieval system. The retrieval system poses challenges in accurately retrieving relevant information when evaluated either on zero-shot domains Thakur et al. (2021) or low-resource languages Zhang et al. (2023). The incorrect or non-relevant information contained in retrieved passages can mislead the LLM to hallucinate Adlakha et al. (2023); Chen et al. (2023); Shi et al. (2023); Yoran et al. (2023); Yu et al. (2023). However, prior work lacks a comprehensive evaluation of LLM reasoning capabilities in multiple languages. As a result, it is unknown whether the LLM and its utilization of the external information is robust against first-stage retrieval errors across both high- or low-resource languages.
In this work, we introduce NoMIRACL, a new dataset to evaluate LLM robustness against errors in first-stage external information, i.e., retrieved passages, across 18 typologically diverse languages. We hire over 30 native speakers as annotators Zhang et al. (2023) to construct the dataset. NoMIRACL contains two subsets, non-relevant and relevant. Where, the non-relevant subset contains queries with no-known answers, i.e., all top-$k$ retrieved passages manually judged to non-relevant. Conversely, the relevant subset contains queries with known answers, i.e., at least one of the top-$k$ passages is manually judged relevant.
We build an initial GPT-4 baseline Brown et al. (2020), a closed-book multilingual LLM on NoMIRACL using the top-$k$ oracle passages retrieved using a hybrid retrieval system from a language-specific Wikipedia corpus Zhang et al. (2023). We use a vanilla prompt template for GPT-4. Our results show that GPT-4 frequently tends to confuse non-relevant passages as relevant and achieves a high 33.2% hallucination rate on average on the non-relevant subset. On the other hand, GPT-4 achieves a lower 14.2% error rate on average on the relevant subset indicating a strong ability to identify relevant passages with known answers.
Furthermore, we observe a positive correlation (Spearman $\rho$ = 0.39) between the GPT-4 hallucination rate and the language resource size, i.e., Wikipedia corpus size Zhang et al. (2023). We suspect the information required to answer a query is better available in a high-resource language (i.e., harder negative), which potentially confuses the LLM to judge the passage as relevant. These results highlight the need to explore better techniques in future to improve LLM robustness.
Figure 2: Contingency table for robustness evaluation with NoMIRACL. More details are provided in (§2.1); (Subset) denotes the ground-truth in NoMIRACL; (Pred.) denotes the LLM output prediction.
To summarize our contributions are: (i) We introduce NoMIRACL, a novel multilingual dataset to assess LLM robustness against first-stage retrieval errors. (ii) We provide GPT-4 as a baseline and observe challenges in over-reliance on noisy passages in the non-relevant subset over recognizing relevant passages in the relevant subset; We hope NoMIRACL can serve as a valuable dataset towards a much-needed LLM robustness evaluation across diverse languages in RAG.
Figure 3: An overview of the data construction procedure (for English) involved in NoMIRACL.
2 Background and Problem Identification
A challenging issue in RAG is to provide robust and reliable LLM-generated output against first-stage information retrieval system. Overreliance of the LLM on the content of the retrieved passages (i.e., tendency to extract information from passages) can be limiting when passages are noisy or non-relevant Yu et al. (2023); Yoran et al. (2023); Shi et al. (2023).
RAG Background. Retrieval-augmented generation (RAG) (Lewis et al., 2020; Guu et al., 2020) involves a two-stage inference pipeline. In the first stage, given the retrieval system and the user query, the retrieval system provides the subset of top-$k$ passages retrieved from an external data corpus $C$. For the next stage, the user query along with the top-$k$ retrieved passages are provided to the LLM which generates an output summarizing the answer for the query, citing relevant passages in its output.
2.1 Robustness Evaluation
We conduct our robustness evaluation as a contingency table to robustly evaluate the LLM behavior in different scenarios. We evaluate a binary classification task, by comparing LLM predictions against the ground truth provided in NoMIRACL. We provide our contingency table in Figure 2.
Definitions. NoMIRACL contains two subsets, we denote them as either non-relevant ($\mathrm{F}$) or relevant ($\mathrm{T}$). The non-relevant subset contains queries with no-known answers, i.e., all top-$k$ passages are non-relevant, whereas the relevant subset contains queries with known answers, i.e., at least one of the top-$k$ passages is relevant. Similarly, we denote the LLM prediction as either positive ($\mathrm{P}$) indicating the model finds the passage relevant to answer the query and similarly negative ($\mathrm{N}$) denotes the model does not find any passage relevant (i.e., containing the answer) for the query.
Confusion Matrix. In our confusion matrix (cf. Figure 2), False Negative ($\mathrm{FN}$) denotes when the model correctly predicts queries with no-known answers using the noisy retrieved passages, whereas False Positive ($\mathrm{FP}$) denotes when the model prediction is incorrect on the non-relevant subset. Similarly, True Positive (TP) denotes when LLM correctly predicts queries with known answers within the retrieved passages, whereas False Positive (FP) denotes when the model prediction is incorrect (model finds no answer) on the relevant subset.
Metrics. Following prior works Adlakha et al. (2023); Chen et al. (2023), we assess LLM robustness as a binary classification task using two metrics: (i) hallucination rate and (ii) error rate. First, we compute the hallucination rate (in %) $=\mathrm{FP}/\mathrm{FP}+\mathrm{FN}$ which measures the model tendency to hallucinate an answer, when no answer is available in all of the passages in the non-relevant subset. Next, we measure the error rate (in %) $=\mathrm{TN}/\mathrm{TN}+\mathrm{TP}$ which measures the model ability to identify the relevant passages in the relevant subset.
Split / ISO ar bn de en es fa fi fr hi id ja ko ru sw te th yo zh Total
Non-relevant Subset: Queries with all human-judged non-relevant passages
Development 228 495 171 289 245 760 98 1,016 1,016 474 211 1,577 268 508 480 323 1,678 1,085 10,922
Test 291 630 218 367 311 968 125 1,294 1294 603 269 2,006 342 646 610 412 2,136 1,381 13,903
Relevant Subset: Queries with at least one human-judged relevant passage
Development 2,896 411 305 799 648 632 1,271 343 350 960 860 213 1,252 482 828 733 119 393 13,495
Test 1,405 1,130 712 1,790 1,515 1,476 801 711 819 611 1,141 1,417 718 465 793 650 663 920 17,737
Table 1: Dataset Statistics for NoMIRACL. The dataset contains two subsets for all 18 languages: (i) Non-relevant subset, where queries contain all human-judged non-relevant passages. (ii) Relevant subset, where queries contain at least one relevant human-judged passage. Both subsets are split into disjoint development and test splits.
3 NoMIRACL Dataset
In this section, we describe our NoMIRACL construction procedure in (§3.1), fold creation in (§3.2), difficulty in providing answers for queries in NoMIRACL in (§3.3), and languages covered and dataset usage in (§3.4). An overview of our data construction procedure can be found in Figure 3.
3.1 Data Construction Procedure
NoMIRACL, a supplementary dataset, was created in the course of developing MIRACL Zhang et al. (2023). The data annotation procedure is conducted in two stages, following Zhang et al. (2023); Clark et al. (2020). In the first stage, the annotator (a native language speaker) is asked to generate a well-formed query for each individual prompt text. Each prompt is a short text snippet containing the first 100 words from a language-specific Wikipedia corpus $C$. Next, for each human-generated query, top-$k$ passages are retrieved from the Wikipedia corpus $C$ using a hybrid multilingual retrieval system (more details in §4.1). In the next stage, annotators assess the relevance of the top-$k$ query–passage pairs as a binary judgment, either relevant (relevance = $1$) or non-relevant (relevance = 00). For more details regarding the annotation procedure, such as the hiring process or stipend, we would like to refer the reader to Zhang et al. (2023).
Non-relevant subset. Annotators generate queries based on certain dataset guidelines, however, occasionally the generated queries cannot be answered with external corpus, which leads to the scenario where none of the top-$k$ passages are relevant, i.e., contains the answer. These queries with no-known answers occur during an annotation procedure due to the following reasons: (i) Queries can be either generic or specific for information to be present in Wikipedia, for e.g. “What does the AC button on a calculator stand for?” retrieves the Wikipedia page on Calculator, [^2] https://en.wikipedia.org/wiki/Calculator but it does not contain information about the AC button; (ii) Spelling mistakes in query generation. We construct the NoMIRACL non-relevant subset with queries with all top-$k$ passages judged as non-relevant by an annotator with a relevance of 00.
Relevant subset. Queries with at least one relevant passage are included in MIRACL Zhang et al. (2023). These queries contain known answers, as at least one of top-$k$ retrieved passages is marked relevant to provide sufficient information to answer the query. We construct the NoMIRACL relevant subset with queries with at least one relevant passage, i.e., all query-passage pairs have been explicitly judged by an annotator as either relevant with a relevance of $1$ or non-relevant with relevance of 00.
3.2 Fold Creation
In NoMIRACL, we split the non-relevant and relevant subsets to form disjoint development and test splits. Detailed statistics can be found in Table 1.
Development split. For queries present in the relevant subset, we reuse the queries from the MIRACL development split Zhang et al. (2023) for all 18 languages. For queries in the non-relevant subset, we randomly sample a disjoint set containing 44% of the queries from the whole non-relevant subset.
Test split. For queries present in the relevant subset, we reuse the queries from the MIRACL test-B split Zhang et al. (2023) for all 18 languages. [^3] We left out the MIRACL test-A split, as it contains queries for only 10 out of the 18 languages available. For queries in the non-relevant subset, we utilize the other disjoint set, containing 56% of the queries from the whole non-relevant subset.
3.3 No Available Answers
Coming up with reliable and accurate answers for a given query requires domain experts as annotators. Annotators can potentially highlight short extractive spans of answers within relevant passages, however, non-extractive queries can either contain multiple answers or a long-form answer, making it difficult to highlight a relevant answer span. Therefore, for NoMIRACL, we focus on the evaluation of top-$k$ passages as information contexts, which were easily judged for their relevancy by a data annotator.
3.4 Languages Covered and Dataset Usage
NoMIRACL covers 18 diverse typological languages Zhang et al. (2023). We cover 10 language families (from Niger-Congo to Indo European), 11 scripts (from Latin to Devanagari) covering diversity from the perspective of linguistic characteristics. The languages along with their ISO codes are listed below: Arabic (ar), Bengali (bn), German (de), English (en), Spanish (es), Persian (fa), Finnish (fi), French (fr), Hindi (hi), Indonesian (id), Japanese (ja), Korean (ko), Russian (ru), Swahili (sw), Thai (th), Yoruba (yo), Chinese (zh).
NoMIRACL is not a static dataset. From Table 1, we observe an uneven amount of queries present in each language. To avoid this, at runtime, a user provides a sampling ratio for queries in the non-relevant to relevant subset to utilize the maximum amount of queries available in NoMIRACL. Using the sampling ratio, we compare each subset (non-relevant and relevant) and return the maximum amount of queries possible.
4 Experimental Setup
Question-answering datasets, such as Natural Questions Kwiatkowski et al. (2019) or WebQuestions Berant et al. (2013) include a well-defined extractive answer for a majority of queries present within the dataset. In contrast, NoMIRACL does not contain answers as highlighted in (§3.3), but rather human-judged relevance judgements for all top-$k$ retrieved passages for 18 different languages. Therefore, we wish to evaluate LLM robustness, i.e., whether the query can be answered using the retrieved passages in RAG.
4.1 Evaluation Setup and Metrics
In NoMIRACL, we evaluate LLM hallucination using an input query, a vanilla prompting technique and top-$k$ (oracle) retrieved passages.
Retrieved Passages. For each query in NoMIRACL, a maximum of $k=10$ passages are retrieved and judged by our annotators. We follow the hybrid retrieval setup in Zhang et al. (2023), which includes three different multilingual retriever models: (i) BM25 Robertson and Zaragoza (2009), a lexical retriever, previously shown to be robust across domains and languages Thakur et al. (2021); Zhang et al. (2022). We use the BM25 implementation available in Anserini Yang et al. (2018) with default parameters and the corresponding language-specific analyzer. (ii) mDPR Karpukhin et al. (2020), a dense retriever, using mBERT Devlin et al. (2018) as the backbone and fine-tuned on MS MARCO with the Tevatron toolkit Gao et al. (2023a). (iii) mColBERT Khattab and Zaharia (2020), a multi-vector retriever, fine-tuned following Khattab and Zaharia (2020) using mBERT as backbone and fine-tuned on MS MARCO. The top-$k$ passages are ranked using an ensemble fusion by normalizing each individual model score within the range $[0,1]$ and further averaging the scores.
Robustness Evaluation. In our work, we evaluate LLM robustness as a response string $y$ in a binary classification task. Following prior evaluation strategies in Adlakha et al. (2023); Yu et al. (2023), we use the input query $q_{i}$, a vanilla prompt template $Q$, and a set of top-$k$ annotated passages $P_{k}$. We prompt the LLM to evaluate whether $q_{i}$ can be answered using any passage in $P_{k}$. The LLM generates a binary output containing either $y$ = “I don’t know” as $\mathrm{N}$ or “Yes, answer is present” as $\mathrm{P}$. The output is tagged “Invalid” if it does not fall in either one of the above. Recall from (§2.1), we calculate the hallucination rate (in %) = $\mathrm{FP}/\mathrm{FP}+\mathrm{FN}$, which measures the error in rejecting information from non-relevant passages and the error rate (in %) = $\mathrm{TN}/\mathrm{TN}+\mathrm{TP}$, which measures the error in identifying relevant passages amongst noisy ones.
4.2 Baseline Model: GPT-4
NoMIRACL provides a baseline based on a closed-book LLM: GPT-4 OpenAI (2023). GPT-4 has been trained to follow an instruction via a prompting strategy Ouyang et al. (2022) and provides a detailed response. Our choice lies with GPT-4 for its multilingual capabilities in machine translation Jiao et al. (2023), and natural language inference tasks Ahuja et al. (2023).
Model Settings. For our experiments, we evaluate GPT-4, using the OpenAI service (API version 2023-05-15) deployed on Microsoft Azure.444learn.microsoft.com/en-us/azure/ai-services/openai/ [^5] We compared Azure API with OpenAI API across four languages in NoMIRACL and observed no noticeable difference between different GPT-4 API version providers. For all our experiments, we maintain a context window size of a maximum of 4096 tokens, a low-temperature score = 0.3 for a deterministic LLM output, and top-p sampling ratio = 0.95. We output a maximum of $50$ tokens from GPT-4.
I will give you a question and several contexts containing information about the question. Read the contexts carefully. If any of the contexts answers the question, respond as either ‘‘Yes, answer is present’’ or ‘‘I don’t know’’:
QUESTION: {query}
CONTEXTS:
[1] {Passage title}: {Passage text}
[2] {Passage title}: {Passage text}
[10] {Passage title}: {Passage text}
OUTPUT:
Figure 4: Vanilla zero-shot prompt template used in our work with GPT-4 OpenAI (2023) for all 18 languages in NoMIRACL. The instruction is provided in English, similar to Ahuja et al. (2023).
Figure 5: GPT-4 hallucination rate ($\mathrm{FP}/\mathrm{FP}+\mathrm{FN}$) on the non-relevant subset ($\mathrm{F}$) in NoMIRACL test split. The non-relevant subset contains queries with no-known answers, i.e., all top-$k$ passages are judged by a human annotator as non-relevant. On average, GPT-4 OpenAI (2023) achieves a hallucination rate of 33.2% to dismiss information available in non-relevant (noisy) retrieved passages. Lower the hallucination rate is better.
Figure 6: GPT-4 error rate ($\mathrm{TN}/\mathrm{TN}+\mathrm{TP}$) on the relevant subset ($\mathrm{T}$) in NoMIRACL test split. The relevant subset contains queries with known answers, i.e., at least one of the top-$k$ ($k=10$) passages Zhang et al. (2023) judged by a human annotator is relevant. On average, GPT-4 OpenAI (2023) achieves an error rate of 14.9% to identify relevant information (relevant in answering the query) within the retrieved passages. Lower the error rate is better.
Vanilla Prompting. The choice of prompt significantly influences the performance and LLMs have been shown brittle to prompting variations, training examples or long context setups Liu et al. (2023). In this work, we evaluate GPT-4 using a zero-shot monolingual listwise prompting strategy. We construct a vanilla prompt template using all top-$k$ (oracle) passages (available in NoMIRACL) as a list of contexts along with the input query, both in the same language. We evaluate GPT-4 zero-shot, as we cannot fit few-shot exemplars due to insufficient context length. Our template provides a short description in English describing the task (Ahuja et al., 2023). Our prompt can be found in Figure 4.
Reducing Costs. Running GPT-4 is expensive. For long contexts and low-resource languages, the costs even can multiply. To limit our experimental cost, we did not exceed our prompt above 4096 tokens. To effectively fit all $k=10$ passages within the vanilla zero-shot prompt, we truncate each passage to use the first 390 tokens. Next, for evaluation, we use a sampling ratio of 50-50% (non-relevant to relevant subset) and limit GPT-4 evaluation to a maximum of 250 randomly sampled queries for all languages in both NoMIRACL relevant and non-relevant split. In our experiments, we ended up producing $\approx$10K API calls producing an expense of \$1,174 (in USD) including miscellaneous costs.
5 Experimental Results
We discuss our GPT-4 baseline results using the hallucination rate on the non-relevant subset in (§5.1) and the error rate on the relevant subset in (§5.2).
5.1 NoMIRACL Non-relevant Subset
Figure 5 shows GPT-4 hallucination rates on the NoMIRACL non-relevant subset for a maximum of 250 queries evaluated (each language) on 18 languages. On average, GPT-4 achieves a 33.2% hallucination rate, which highlights the challenge to identify all non-relevant passages in retrieved information in the prompt. The lowest hallucination rates (marked in green) are found in Swahili (sw) with 8.8% and Spanish (es) with 11.6%. Whereas, the highest (marked in red) are found in Thai (th) with 53.2% and French (fr) with 46.8%.
Further, we calculate the correlation of hallucination rates in GPT-4 with resource availability, i.e., Wikipedia corpus size in NoMIRACL. We achieve a positive correlation (Spearman’s rank correlation $\rho=0.39$, which can increase up to $\rho=0.51$ without Spanish). We hypothesize that queries in lower and medium-resource languages (smaller Wikipedia corpus) contain retrieved information that is likely to be non-relevant (easier negative), thereby making it easy for the LLM to judge as “I don’t know”. Future work can focus on improving GPT-4’s ability not to generate an answer for non-relevant knowledge. We also note that our current baseline uses a closed-box or black-box LLM; we encourage future work to explore open-source alternatives on NoMIRACL.
5.2 NoMIRACL Relevant Subset
Figure 6 shows the GPT-4 error rates on NoMIRACL relevant subset for a maximum of 250 queries (each language) on 18 languages. On average, GPT-4 achieves a low error rate of 14.9%, which is 20.3% lower than the hallucination rate, indicating a stronger ability to identify relevant passages in retrieved information. GPT-4 observes the lowest error rates with below 5% for high-resource languages such as Chinese (zh), Thai (th), Japanese (ja) or English (en). The highest error rate is observed in Telugu (te) with 42.0%. Overall, the low error rates indicate that model is able to identify relevant passages present within non-relevant passages, with a relatively higher error rate visible for low-resource languages, for example, Telugu (te), Hindi (hi) and Swahili (sw). Spanish (es) is a unique exception in the high-resource language category, where GPT-4 dismiss information confidently as “I don’t know” with a 22.0% error rate on the relevant subset.
6 Ongoing Work
Identifying LLM reasoning capabilities and robustness against first-stage retrieval errors in RAG applications has only begun. The first step in our robustness evaluation is to release the NoMIRACL dataset with both the development and test splits for all languages and provide a GPT-4 baseline to highlight the challenge in LLM robustness and facilitate more robust and accurate LLM-generated outputs in the future.
We are actively working on benchmarking a wider variety of open-source multilingual LLMs, a few notable ones are LLAMA-2 Touvron et al. (2023), FLAN-T5 Chung et al. (2022) and PolyLM Wei et al. (2023) among others. We also wish to evaluate whether better prompting strategies or fine-tuning LLMs can help improve robustness by evaluating methods such as Chain-of-Thought (CoT; Wei et al., 2022), Chain-of-Verification (CoVe; Dhuliawala et al., 2023) or recently introduced Chain-of-Noting (CoN; Yu et al., 2023).
7 Related Work
Retrieval-Augmented Generation. The knowledge stored in a large language model (LLM) is commonly outdated He et al. (2023), and prone to hallucinations by generating factually incorrect output Maynez et al. (2020); Raunak et al. (2021). By grounding on external knowledge, a retrieval-augmented LLM can generate better and more trustworthy output Guu et al. (2020); Lewis et al. (2020); Izacard and Grave (2021); Borgeaud et al. (2022). Retrieval-augmented generation has achieved remarkable results in various tasks such as open-domain question answering (ODQA) Lewis et al. (2020); Izacard and Grave (2021); Trivedi et al. (2023), argument extraction Du and Ji (2022) and code generation Zhou et al. (2023). Real-world products such as Bing Search [^6] New Bing Search: news.microsoft.com/the-new-Bing/ and Langchain [^7] Langchain: github.com/langchain-ai/langchain have incorporated RAG applications.
LLM Evaluation. Prior work explores adding perturbation in passages and shows that LLM performance can be influenced when exposed to different tasks, such as question answering (QA)  Jia and Liang (2017); Petroni et al. (2020); Creswell et al. (2023), logical reasoning  Misra et al. (2023) or arithmetic reasoning  Shi et al. (2023); Kumar et al. (2021). In examining controllability and robustness, Li et al. (2023) observes that LLMs disregard contextual information, showing that LLM output can be influenced by non-relevant context.  Adlakha et al. (2023) observes complementary results from our work, where they observe LLM can be rather faithful when provided non-relevant passages in QA datasets such as NQ Kwiatkowski et al. (2019). Knowing that prompting LLMs with non-relevant data can result in misguided responses, Yu et al. (2023) recently introduced a new prompting technique, Chain-of-Noting (CON) and Yoran et al. (2023) fine-tuned the LLM explicitly, both aimed to improve LLM robustness in RAG when non-relevant information is provided.
Related Datasets. Datasets focused on addressing unanswerable queries such as SQuAD 2.0 Rajpurkar et al. (2018) were created in an adversarial way to look similar to datasets with answerable queries. Similarly, Conversational QA datasets such as CoQA Reddy et al. (2019) and QuAC Choi et al. (2018) also contain unanswerable queries. A concurrent work proposes RGB, a RAG benchmark to evaluate LLM robustness in English and Chinese Chen et al. (2023).
8 Conclusion
Retrieval-augmented generation setups are effective in the factual grounding of LLM-generated output with information available in external knowledge bases. However, in cases where retrieved information does not assist in providing the answer, a robust LLM should potentially identify them. In our work, we provide NoMIRACL for evaluating LLM robustness as a binary classification task on 18 languages. We provide two subsets in NoMIRACL, the non-relevant subset, where queries contain all non-relevant passages, and the relevant subset, where queries contain at least a single relevant passage. We evaluate robustness using two metrics: hallucination rate and error rate.
In our work, we build a GPT-4 baseline on NoMIRACL. GPT-4 achieves a high 33.2% hallucination rate on the non-relevant subset and 14.9% error rate on the relevant NoMIRACL split, highlighting that GPT-4 finds it challenging to dismiss non-relevant passages over relevant passages in first-stage retrieved information. We open-source NoMIRACL and hope it encourages future work to improve LLM robustness.
Acknowledgements
We would like to thank our annotators, without whom NoMIRACL could not have been built. This research was supported in part by the Natural Sciences and Engineering Research Council (NSERC) of Canada, a gift from Huawei, and Cloud TPU support from Google’s TPU Research Cloud (TRC).